Use the fact that 0-a= -a to justify that
-(-5) equals 5.
1
Expert's answer
2020-10-27T20:35:25-0400

We have given the fact that

"0+(-a)=-a"

Claim:"-(-a)=a"

Proof: Note that We can write "0=a+(-a)=(-a)+a"

Now add additive inverse of "(-a)" to the right, which is "-(-a)" to the original equation, we get


"0+(-a)+(-(-a))=-a+(-(-a))\\
\iff 0+0=-a+(-(-a))\iff0=-a+(-(-a))\\"

Now add "a" to the left in above last equation, thus


"a+0=a=a+(-a)+(-(-a))\\
\iff a=0+(-(-a))=(-(-a))=-(-a)"

Proved.


Now put "a=5" in the above claim, we get the desired result.